OpenStudy (anonymous):

please explain thi i dont get it:( Which of the following are the x and y intercepts of the equation 4x - 2y = 8?

OpenStudy (aripotta):

change it to slope-intercept form first

OpenStudy (anonymous):

-2y=-4x+8?

OpenStudy (aripotta):

divide everything by -2, then you'll be in slope-intercept form

OpenStudy (anonymous):

y=2x+-4

OpenStudy (aripotta):

y = 2x - 4*

OpenStudy (aripotta):

so now you can tell what the y intercept is. what is it?

OpenStudy (anonymous):

what is a y intercept?

OpenStudy (aripotta):

:l

OpenStudy (anonymous):

Basically, what you do is you change it to y=mx+c first. If we rearrange it like that, it'd look like this: \[y=2x-4\] Then you know that the y-intercept is -4. To find the x-intercept, you make y=0 and solve for x. \[0=2x-4\] \[4=2x\] \[x=2\] So the y-intercept is -4 and the x-intercept is 2.

OpenStudy (anonymous):

ok what makes it a y intercept?

OpenStudy (anonymous):

We know that when y=mx+c. C is always the y-intercept and m is the slope :)

OpenStudy (anonymous):

ok so the answer is the last one?(3, 0) and (0, 4) (2, 0) and (0, 4) (-2, 0) and (0, -4) (2, 0) and (0, -4)

OpenStudy (anonymous):

Yes I would assume so!

OpenStudy (anonymous):

i have a computer block thing i cant watch videos sorry=(